OpenStudy (kaley246):

Which is the set of roots of the equation? 0 = x2 + 10x - 24 (Points : 4) {12, 2} {-12, -2} {-12, 2} {12, -2}

OpenStudy (kaley246):

@SolomonZelman

OpenStudy (solomonzelman):

\(\large\color{green}{ \displaystyle 24=12\times(-2) }\) \(\large\color{blue}{ \displaystyle 10=(-2)+(12) }\)

OpenStudy (kaley246):

so my answer is -2,12

OpenStudy (solomonzelman):

when you factor you get: \(\large\color{black}{ \displaystyle x^2+10x-24~~~\Rightarrow~~~(x-\color{blue}{\bf 2})(x+\color{blue}{\bf 12}) }\)

OpenStudy (solomonzelman):

(( either `(x-2)=0`, or `(x+12)=0` ))

OpenStudy (solomonzelman):

you are a little off.

OpenStudy (kaley246):

whats the answer then

OpenStudy (solomonzelman):

Do you understand how I factored the equation ?

OpenStudy (kaley246):

The graph represents the height of a football t seconds after it was punted. How many seconds did it take the football to land? (Points : 4) 2 5 3 4

OpenStudy (solomonzelman):

we haven't got through the previous question.... have we ?

OpenStudy (solomonzelman):

and for this question there is an Info lacking.